OpenStudy (anonymous):

What is the vertex of the parabola whose equation is y = (x- 1)2 + 3? (-1, -3) (-1, 3) (1, 3) (1, -3)

OpenStudy (anonymous):

the vertex has x coordinate 1 ( from (x-1) at x = 1 y = (1-1)^2 + 3 = 3 no 3 is correct

OpenStudy (zepp):

This is the standard form of a quadratic function, f(x) = a(x-h)^2 + k Where (h,k) is the vertex

OpenStudy (zepp):

In this case, y = (x-1)^2 + 3 (1,3) is the vertex. Just some reminder: If your equation was y = (x + 1)^2 + 3 (Notice the plus sign) (-1,3) would be the vertex.